OpenStudy (anonymous):

Solve: |2n - 3| = 5 {-1, 4} {-1, 1} {1, -4}

OpenStudy (angela210793):

Halo what don't you try solving this yourself first???

OpenStudy (anonymous):

because i dont understand how to do these...

hero (hero):

In general: |ax + b| = c becomes: ax + b = -c and ax + b = c In this case, |2n - 3| = 5 becomes: 2n - 3 = -5 and 2n - 3 = 5 Solve for n: 2n = -5 + 3 2n = 5 + 3 2n = -2 2n = 8 n = -2/2 n = 8/2 n = -1 n = 4

OpenStudy (anonymous):

Absolute values always have two cases. The positive and negative cases here are\[2n - 3 = 5\rightarrow 2n=8\rightarrow n=4\]\[2n-3=-5\rightarrow 2n=-2\rightarrow n=-1\]So your two solutions are n=-1 and n=4

OpenStudy (anonymous):

I guess I need to type faster.

hero (hero):

Why?

OpenStudy (anonymous):

thank you both :) lol couldu guys possibly help with my next question.. its the last one i reall am not understanding thesee.. algebra hates me D: